OpenStudy (anonymous):

\[\int\limits_{3}^{3x^4} t sint^4 dt\]

OpenStudy (anonymous):

use integral by parts and apply the limits

OpenStudy (anonymous):

how do you do that?

OpenStudy (anonymous):

take u =t^4 and v= sin u

sam (.sam.):

\[\int\limits t \sin \left(t^4\right) \, dt \]or \[\int\limits t \sin ^4(t) \, dt\]